the length of a new rectangular is 3 yards longer then quadruple the width. if the perimeter is 606 what is the dimensions.
Answer by algebrahouse.com(1659)

x = width
4x + 3 = length {three yards longer than quadruple the width}

Perimeter of a rectangle = 2(width) + 2(length)

2x + 2(4x + 3) = 606 {substituted width and length into perimeter formula}
2x + 8x + 6 = 606 {used distributive property}
10x + 6 = 606 {combined like terms}
10x = 600 {subtracted 6 from each side}
x = 60 {divided each side by 10}
4x + 3 = 243 {substituted 60, in for x, into 4x + 3}

width = 60 yds
length = 243 yds

For more help from me, visit: www.algebrahouse.com